OpenStudy (help!!!!):

A phonograph record turns at the rate of 33 1/3 revolutions per minute. In how many seconds will the record revolve 180 Degrees?

OpenStudy (help!!!!):

I know that 33 1/3 can be 100 revolutions and I know that a revolution is equal to 360 degrees and a minute has 60 secs. How do I solve for this?

OpenStudy (aravindg):

First you need to convert revolution into degrees and minute in seconds. 1 Revolution=360 degree 1 Minute=60 seconds

OpenStudy (help!!!!):

yup

OpenStudy (help!!!!):

So would there be 50 revolutions per minute? Thus there are 6/5 per 180 degrees? @AravindG

OpenStudy (help!!!!):

So there are 50 180degrees per minute right? So there are 5/6 seconds in 180?

OpenStudy (aravindg):

33 1/3=100/3 Right? You multiply that with 360 degrees to find angle turned.

OpenStudy (help!!!!):

so 12000? What is next?

OpenStudy (help!!!!):

@TheSmartOne @zepdrix

OpenStudy (aravindg):

Yep.

OpenStudy (aravindg):

Now convert minute to seconds and divide the angle by time in seconds.

OpenStudy (help!!!!):

@TheSmartOne @zepdrix

OpenStudy (help!!!!):

Cool so 72000 seconds is the conversion. Then do this 180/72000 or 360/72000? @AravindG

OpenStudy (help!!!!):

Here are the answer choices btw 1/3, 3, 1 4/5, 9/10, 10

OpenStudy (aravindg):

What dod you get when you divided the angle and time in seconds?

OpenStudy (help!!!!):

Which angle are you talking about 360 or 180?

Directrix (directrix):

When the record spins exactly one time back around to its starting position, a point on the rim of the record has traveled 360 degrees. In 33 1/3 revolutions, a point on the rim of the circle travels 33 1/3 times 360 degrees per minute. A minute is 60 seconds. 33 1/3 = 100/3 as an improper fraction In 100/3 revolutions, a point on the rim of the circle travels 100/3 times 360 degrees per 60 seconds. In how many seconds will the record revolve 180 Degrees? Set up a proportion: (100/3) * 360 180 ------------ = ----- 60 x where x is the number of seconds for a 180 degree spin. 12000 = 180 ----- ----- 60 x x = ? @HELP!!!!

Directrix (directrix):

@HELP!!!! Any questions?

OpenStudy (help!!!!):

Thanks!!

Directrix (directrix):

You are welcome. I hope the proportion makes sense.

OpenStudy (help!!!!):

Yup I does! Thanks for the clarification!!